day shift, 8 July 99

Marj Corcoran (corcoran@physics.rice.edu)
Thu, 8 Jul 1999 15:56:50 -0500 (CDT)


Shift Summary for day shift, 8 July 99

On Shift: M. Corcoran and V. Prasad, with support from C. Wolfe
P. Toale

Hours of beam available at all intensities: 3.0
Hours of beam available at nominal intensity: 3.0
Hours of E832 data (special runs not included): 3.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: rad trip, MI RF problems, TeV quench
========================================================================

Runs ended this shift:
Run 13837 E832 data
Run 13838 in progress

=======================================================================

Tapes brought to FCC this shift: QKC095-097
QKS095-097
QKK095-097
QKB033

========================================================================
Accesses during shift: None
========================================================================

Description of Shift: Fairly smooth data-taking while there was beam.

SEM was recalibrated during run 13838; all actual data in this run
occurred after the recalibration. But we were totally confused...the
SEM value on the TV screen does not agree with the SEM coming from the
slow daq; for example TV says 4.16E12 while slow daq says 3.45E12.
Evening shift will try to figure out what's going on.

New He trailer installed; flow rates increased slightly on DC1 and DC4.

======================================================================
run spill time comment
------------------------------------------------------------------------
13837 7:15 Day shift takes over, run in progress
13837 250 7:30 Beam gone, KTeV rad trip for unknown reasons
13837 260 7:40 Beam returns
13837 290 8:15 Beam gone again, TeV power lead problems and
MI RF problems, est 1 hr
13837 366 9:55 Beam returns
13837 372 10:05 HCC inefficient
13837 385 10:18 Etot reloaded, HCC now OK. Spills 372-385 had
inefficient HCC
13837 467 12:10 Tapes fill, run ends, tapes swapped and
new run started
13838 1 12:20 New run starts and TeV immediately quenches. No
actual data on tape before quench
13838 15:00 Rick Ford changed the calibration of the SEM, it
now reads lower by 21%, the request has been
lowered to 4E12 to maintain the same true intensity.
This change was made before any actual beam in
this run.
13838 119 15:10 Beam finally returns
========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

Run a looong laser scan if we have at least 1 hr down time.
Clear up confusion about SEM.